# MSTN: Fast and Efficient Multivariate Time Series Model

### v1

@@ Abstract
Real-world time-series data is highly non stationary and complex in dynamics that operate across multiple timescales, ranging from fast, short-term changes to slow, long-term trends. Most existing models rely on fixed-scale structural priors, such as patch-based tokenization, fixed frequency transformations, or frozen backbone architectures. This often leads to over-regularization of temporal dynamics, which limits their ability to adaptively model the full spectrum of temporal variations and impairs their performance on unpredictable, Sudden, high-magnitude events. To address this, we introduce the Multi-scale Temporal Network (MSTN), a novel deep learning architecture founded on a hierarchical multi-scale and sequence modeling principle. The MSTN framework integrates three core components: (i) a multi-scale convolutional encoder that constructs a hierarchical feature pyramid for local patterns (ii) a sequence modeling component for long-range temporal dependencies. We empirically validate this with Bidirectional Long Short-Term Memory (LSTM) and Transformer variants, establishing a flexible foundation for future architectural advancements. and (iii) a gated fusion mechanism augmented with squeeze-and-excitation (SE) and multi-head temporal attention (MHTA) for dynamic, context-aware feature integration. This design enables MSTN to adaptively model temporal patterns from milliseconds to long-range dependencies within a unified framework. Extensive evaluations across time-series long-horizon forecasting, imputation, classification and generalizability study demonstrate that MSTN achieves competitive state-of-the-art (SOTA) performance, showing improvements over contemporary approaches including EMTSF, LLM4TS, HiMTM, TIME-LLM, MTST, SOFTS, iTransformer, TimesNet, and PatchTST. In total, MSTN establishes new SOTA performance on 24 of 32 benchmark datasets, demonstrating its consistent performance across diverse temporal tasks.
@@
[…]
[inst2]organization=Department of Civil Engineering, Indian Institute of Technology Indore, addressline=Indore, postcode=452020, state=Madhya Pradesh, country=India
@@ 1 Introduction
Multivariate time series analysis is a crucial component of modern AI, with significant real-world applications ranging from healthcare monitoring and industrial predictive maintenance to behavioral analytics. It has been extensively studied for many years and has been widely used in tasks such as weather forecasting, imputing missing data for data mining, and classification of trajectories for action recognition [wu2022autoformer, zhao2024himtmhierarchicalmultiscalemasked]. Due to its significant practical value, time series analysis has attracted considerable attention. Unlike language or video, time-series data consists of individual scalar measurements, where single time points rarely carry sufficient semantic meaning. This fundamental characteristic forces models to extract information from temporal variation patterns of continuity, periodicity, and trend, making robust temporal modeling exceptionally challenging [wu2023timesnettemporal2dvariationmodeling, lim2021time].
@@
Deep learning (DL) models have demonstrated strong performance not only in forecasting tasks, but also in representation learning, where abstract features can be extracted and transferred to downstream tasks such as classification and forecasting, achieving SOTA results [nie2023timeseriesworth64]. Recent findings show that simple linear models can outperform sophisticated Transformer variants on common benchmarks [zeng2022transformerseffectivetimeseries], highlighting fundamental limitations in contemporary architectures. This is particularly critical in safety-sensitive domains like behavioral analytics and industrial monitoring, where patterns are often non-periodic and evolve across multiple timescales. The design of temporal models faces the constant challenge of balancing model strength with computational efficiency. Models like recurrent neural networks (RNN) and their gated variants [6795963, lai2018modelinglongshorttermtemporal] offer a strong sequential modeling but suffer from vanishing gradients and limited parallelism. Temporal convolutional networks (TCN)  [he2019temporal] improve computational efficiency but are constrained by fixed receptive fields. The Transformer architecture  [vaswani2023attentionneed] originally promised global dependency modeling, but its quadratic complexity prompted a line of research focused primarily on sparsification. Models like Informer [zhou2021informer] and FEDformer [zhou2022fedformer] develop efficient attention variants, yet often retain a point-wise view of time series, treating each time step as an independent token and struggling to capture local semantic structures.
These challenges have shaped several predominant, yet ultimately constrained, modern paradigms. One approach, exemplified by PatchTST [nie2023timeseriesworth64], adapts the patching technique from Vision Transformers, grouping adjacent time points into sub-series tokens. This approach successfully captures local semantics and benefits from channel-independent processing, but its reliance on a fixed patch size creates a fundamental rigidity, locking the model into a single, pre-defined scale of analysis. Another paradigm, improved by TimesNet [wu2023timesnettemporal2dvariationmodeling], reframes the problem by transforming 1D time series into 2D tensors based on identified periods. While powerful for data with strong periodic components, this approach is intrinsically dependent on reliable periodicity detection and the 2D transformation prior, making it less suited for non-stationary, irregular, or non-periodic patterns where such clear structure is absent. Building on these concepts, MTST [zhang2024multiresolutiontimeseriestransformerlongterm] employs a multi-resolution patch-based design with a multi-branch architecture for simultaneous modeling of diverse temporal patterns at different resolutions. However, its fixed choice of resolutions is less flexible and may fail to generalize across datasets with irregular or heterogeneous temporal patterns. Patch tokenization can also lose fine-grained dynamics, while the reliance on relative positional encoding restricts robustness in non-seasonal or event-driven sequences.
We identify a fundamental limitation on architectures that impose specific single-scale structural constraints, such as a fixed patch size, a period-based 2D shape, or frozen backbone architectures. These prior models struggle with the inherently multi-scale nature of many real-world temporal incidents. Their architecture limits their ability to model discriminative patterns that manifest simultaneously across different timescales, from brief micro-events to slow macro and long-term trends, such as found in risky driving behaviour. This gap highlights the need for a more flexible temporal representation that can primarily adapt to this hierarchical structure without relying on a single, potentially misaligned, structural assumption.
To address these challenges, we propose a DL model (MSTN) that naturally captures temporal hierarchies without relying on fixed and repeated structural assumptions. MSTN unifies multi-scale feature learning with global sequence modeling through a dual-path architecture with a sequence modeling core. We empirically validate this design with two instantiations: i) MSTN–BiLSTM, which excels at modeling local continuity and behavioral smoothness, and ii) MSTN–Transformer, which captures local complex patterns and long-range contextual dependencies. In both variants, convolutional encoders extract fine-grained local features, while the temporal core captures long-range dependencies. Their outputs are integrated via a gated fusion mechanism enhanced with SE recalibration and MHTA, yielding compact, discriminative, and temporally coherent representations. This design provides an unprecedented trade-off between representational capacity and computational efficiency.
@@
We demonstrate that MSTN establishes new SOTA performance on 24 of 32 benchmark datasets for time series analysis. This includes 9 forecasting benchmarks, and 6 imputation tasks and 17 classification datasets (10 from the standard UEA archive and 7 additional international benchmarks), demonstrating its versatility and improved performance across diverse applications. MSTN establishes a new SOTA in long-term forecasting, attaining the lowest prediction error on 6 out of 9 standard benchmarks and consistently exceeds leading architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], GPT2(6) [zhou2023fitsallpowergeneraltime], SOFTS [han2024softsefficientmultivariatetime] iTransformer [liu2024itransformerinvertedtransformerseffective], and PatchTST [nie2023timeseriesworth64]. Complementing these results, the MSTN-Transformer variant achieves first-place rankings 31 out of 48 ($\approx$65%) in all 48 imputation scenarios across the ETT, Electricity, and Weather datasets, demonstrating strong resilience under up to 50% missingness. evaluation scenarios and secure the highest average score on 5 out of 6 datasets. In the classification task, MSTN achieves improved performance over recent SOTA architectures such as GTP2(6) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], LigthtTS  [zhang2022morefastmultivariatetime] and FEDformer [zhou2022fedformer]. We further evaluate MSTN for generalizability across 7 benchmark datasets from various domains like human safety, physical activity recognition, animal-welfare monitoring, and mechanical prognostics. Without domain-specific tuning, surprisingly, MSTN achieves uniformly strong performance, ranks first across all datasets, establishing new benchmarks and illustrating its robustness under diverse temporal dynamics. Finally, MSTN-Transformer achieves unprecedented performance with 0.020 MSE, 2.87 MB footprint and 0.72 ms inference time, representing a 22.4× accuracy improvement and 34.7× speedup over SOFTS [han2024softsefficientmultivariatetime]. The model further demonstrates improved classification accuracy (99.53%) with ultra-low latency (0.155 ms) significantly outperforming TimesNet (99.18%, 5.07 ms) and PatchTST (98.64%, 4.80 ms) on the Rodegast et al. [DARUS33012023] dataset. While maintaining compact model sizes (0.54–7.14 MB), this balance of SOTA accuracy, robustness, and efficiency positions MSTN-Transformer as an ideal solution for real-time and edge-AI temporal modeling applications. Our comprehensive evaluation establishes MSTN as a new benchmark, achieving SOTA performance on 24 of 32 datasets across forecasting, imputation, and classification tasks.
@@ 2 Results
This section presents a comprehensive evaluation of the proposed MSTN. Demonstrating its SOTA performance in the long-term forecasting, imputation, classification, and generalizability study. MSTN (MSTN-Transformer and MSTN-BiLSTM) achieves competitive performance, showing improvements over recent DL, Transformer based, and time series approaches.
@@ 2.1 Long-Term Forecasting
Table 1: Multivariate forecasting results with prediction lengths $H\in\{96,192,336,720\}$ for others and for ILI $H\in\{24,36,48,60\}$. Red/Blue: First/Second ranks.
@@
We evaluate long-term forecasting performance using established protocols from the prior literature [wu2022autoformer, wu2023timesnettemporal2dvariationmodeling, nie2023timeseriesworth64] across nine diverse benchmark datasets. The prediction horizons are set to $H\in\{96,192,336,720\}$ for the standard benchmarks and $H\in\{24,36,48,60\}$ for the ILI data set. As presented in Table 1, the MSTN-Transformer and MSTN BiLSTM shows improved performance compared to specialized time series forecasting architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm],SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].
@@
2.2 Imputation
Table 2: Imputation Task Results Comparison. We randomly mask 12.5%, 25%, 37.5%, and 50% time points to compare the model performance under different missing degrees. Red/Blue: First/Second ranks.
(Ours) (Ours) (2023) (2023) (2023) (2022) (2023) (2022)
Mask R. MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
ETTh2
ETTm1
ETTm2
ECL
Weather

### v2

# MSTN: Fast and Efficient Multivariate Time Series Prediction Model

@@ Abstract
Real-world time series often exhibit strong non-stationarity, complex nonlinear dynamics, and behaviour expressed across multiple temporal scales, from rapid local fluctuations to slow-evolving long-range trends. However, many contemporary architectures impose rigid, fixed-scale structural priors—such as patch-based tokenization, predefined receptive fields, or frozen backbone encoders—which can over-regularize temporal dynamics and limit adaptability to abrupt high-magnitude events. To handle this, we introduce the Multi-scale Temporal Network (MSTN), a hybrid neural architecture grounded in an Early Temporal Aggregation principle. MSTN integrates three complementary components: (i) a multi-scale convolutional encoder that captures fine-grained local structure; (ii) a sequence modeling module that learns long-range dependencies through either recurrent or attention-based mechanisms; and (iii) a self-gated fusion stage incorporating squeeze–excitation and multi-head attention to dynamically modulate cross-scale representations. This design enables MSTN to flexibly model temporal patterns spanning milliseconds to extended horizons, while avoiding the computational burden typically associated with long-context models. Across extensive benchmarks covering forecasting, imputation, classification, and cross-dataset generalization, MSTN consistently delivers state-of-the-art performance, outperforming recent leading approaches including TIME-LLM, HiMTM, SOFTS, LLM4TS, TimesNet, and PatchTST, and establishing new best results on 24 out of 32 datasets. Despite its strong performance, MSTN remains lightweight and supports fast inference, making it well suited for deployment on edge devices and resource-constrained environments.
@@ 1 Introduction
Multivariate time series (MTS) analysis has been a classical problem with significant real-world applications including healthcare monitoring [10.1145/3531326], traffic management [kadiyala2014multivariate], environment [gruca2022weather4cast], industrial predictive maintenance, energy  [6714975] and behavioral analytics. It has been extensively studied for many years and has been widely used in tasks such as weather forecasting, imputing missing data for data mining, and classification of trajectories for action recognition [wu2022autoformer, zhao2024himtmhierarchicalmultiscalemasked]. Due to its significant practical value, time series analysis has attracted considerable attention. Unlike language or video, time-series data consists of individual scalar measurements, where single time points rarely carry sufficient semantic meaning. This fundamental characteristic forces models to extract information from temporal variation patterns of continuity, periodicity, and trend, making robust temporal modeling exceptionally challenging [wu2023timesnettemporal2dvariationmodeling, lim2021time].
@@
Deep learning (DL) models have demonstrated strong performance not only in forecasting tasks, but also in representation learning, where abstract features can be extracted and transferred to downstream tasks such as classification and forecasting, achieving state-of-the-art (SOTA) results [nie2023timeseriesworth64]. Recent findings show that simple linear models can outperform sophisticated Transformer variants on common benchmarks [zeng2022transformerseffectivetimeseries], highlighting fundamental limitations in contemporary architectures. The design of temporal models faces the constant challenge of balancing model strength with computational efficiency. Models such as convolutional neural networks (CNNs) [franceschi2019unsupervised, bai2018empirical] and recurrent neural networks (RNNs) with their gated variants [6795963, lai2018modelinglongshorttermtemporal] suffer from vanishing gradients and limited parallelism, which has shifted the paradigm towards more complex models. Temporal convolutional networks (TCN)  [he2019temporal] improve computational efficiency but are constrained by fixed receptive fields. The Transformer architecture  [vaswani2023attentionneed] originally promised global dependency modeling, but its quadratic complexity prompted a line of research focused primarily on sparsification. Models like Informer [zhou2021informer] and FEDformer [zhou2022fedformer] develop efficient attention variants, yet often retain a point-wise view of time series, treating each time step as an independent token and struggling to capture local semantic structures.
These challenges have shaped several predominant, yet ultimately constrained, modern paradigms. These can be formalized into three primary Channel Strategies (CS) [qiu2025comprehensivesurveydeeplearning]: Channel Dependence (CD), which favors high feature completeness at a computational cost; Channel Independence (CI), which prioritizes speed by isolating channels; and Channel Partiality (CP), which aims for a balance by reducing channel complexity. CI based models like PatchTST [nie2023timeseriesworth64] improves local semantic modeling through patch tokenization but introduces rigidity through fixed patch sizes; LLM-derived approaches such as LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] depend on frozen backbone architectures. CD based models such as TimesNet [wu2023timesnettemporal2dvariationmodeling] transforms time series into periodic 2D structures but depends heavily on accurate period discovery, limiting robustness for irregular or non-stationary signals. MTST [zhang2024multiresolutiontimeseriestransformerlongterm] models multiple resolutions but relies on fixed resolution sets and may dismiss fine-grained dynamics. Even highly efficient recent systems such as SOFTS [han2024softsefficientmultivariatetime] prioritize channel efficiency through linear factorization, inherently sacrificing temporal completeness and weakening multi-scale temporal representation. Similarly, the CP Channels strategy in MCformer [10533212] achieves efficiency by partially blending a specific number of channels, but this partiality limits the recovery of full inter-channel context. Consequently, such architectures struggle to sustain fidelity in complex long-sequence environments.
We identify a fundamental limitation on these approaches: the imposition of single-scale architectural assumptions such as fixed patch sizes, fixed periodicity-driven 2D mappings, struggle to capture discriminative cues manifesting across diverse temporal horizons, or static backbone structures. However, real-world temporal phenomena are inherently multi-scale: micro-events, meso-scale fluctuations, and macro trends often coexist. Consequently, existing models struggle to capture discriminative cues manifesting across diverse temporal horizons, particularly in domains such as risky driving behaviour, physiological safety monitoring, or mechanical prognostics—where multi-scale variability is intrinsic. This gap motivates a more flexible, hierarchy-aware temporal representation without relying on a single predefined structure.
To address these limitations, we propose the Multi-scale Temporal Network (MSTN), a hybrid architecture founded on the Early Temporal Aggregation (ETA) principle. MSTN naturally captures hierarchical temporal patterns without fixed structural assumptions through its central innovation: ETA condenses the full temporal input into a single static feature vector via an $L\rightarrow 1$ transformation, effectively neutralizing the $\mathcal{O}(L^{2})$ inference bottleneck while preserving global completeness. This design allows the dual encoder to retain $\mathcal{O}(L^{2})$ modeling capacity during representation extraction, while all subsequent refinement operates in guaranteed constant-time $\mathcal{O}(1)$ complexity. MSTN unifies multi-scale feature extraction with global temporal reasoning through a dual-path architecture: convolutional encoders capture fine-grained local structure, while a Bidirectional Long Short-Term Memory (BiLSTM) or Transformer backbone models long-range dependencies. Their outputs are fused through a self-gated fusion (SGF) mechanism enriched by squeeze-and-excitatio (SE) recalibration and multi-head attention (MHA), producing compact, discriminative, and temporally coherent representations. We empirically validate the MSTN framework through two realizations: MSTN–BiLSTM, ideal for smooth continuous behaviors, and MSTN–Transformer, suitable for complex, long-range event modeling.
@@
We demonstrate that MSTN establishes new SOTA performance on 24 of 32 benchmark datasets for time series analysis. This includes 9 forecasting benchmarks, 6 imputation tasks, and 17 classification datasets (10 from the standard UEA archive and 7 additional international benchmarks), demonstrating its versatility and improved performance across diverse applications. MSTN establishes a new SOTA in long-term forecasting, attaining the lowest prediction error on 6 out of 9 standard benchmarks and consistently exceeds leading architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], GPT2(6) [zhou2023fitsallpowergeneraltime], SOFTS [han2024softsefficientmultivariatetime] iTransformer [liu2024itransformerinvertedtransformerseffective], and PatchTST [nie2023timeseriesworth64]. Complementing these results, the MSTN-Transformer variant achieves first-place rankings 31 out of 48 ($\approx$65%) in all 48 imputation scenarios across the ETT, Electricity, and Weather datasets, demonstrating strong resilience under up to 50% missingness. In the classification task, MSTN achieves improved performance over recent SOTA architectures such as GTP2(6) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], LigthtTS  [zhang2022morefastmultivariatetime] and FEDformer [zhou2022fedformer]. We further evaluate MSTN for generalizability across 7 benchmark datasets from various domains like human safety, physical activity recognition, animal-welfare monitoring, and mechanical prognostics. Without domain-specific tuning, MSTN achieves improved performance, ranking first across all datasets, and illustrates its robustness under diverse temporal dynamics. Concretely, MSTN-Transformer achieves unprecedented performance with 2.87 MB memory footprint, and 0.72 ms inference latency, representing a $22.4\times$ accuracy improvement and $34.7\times$ speedup over SOFTS [han2024softsefficientmultivariatetime]. The model further demonstrates improved classification accuracy (99.53%) with ultra-low latency (0.15 ms), significantly outperforming TimesNet (99.18%, 5.07 ms) and PatchTST (98.64%, 4.80 ms) on the Rodegast et al. [DARUS33012023] dataset. While maintaining compact model sizes (0.54–7.14 MB), this balance of SOTA accuracy, robustness, and efficiency positions MSTN as a preferred solution for real-time and edge-AI temporal modeling applications.
@@ 2 Results
This section presents a comprehensive evaluation of the proposed MSTN model. We demonstrate its performance on three time-series tasks: (1) long-term forecasting, (2) imputation, (3) classification. Additionally, we also present the generalizability study on standard time-series benchmark datasets.
2.1 Benchmark Datasets
Table 1: Benchmark datasets and imputation performance comparison.
(a) Benchmark datasets for forecasting, imputation, and classification (10 UEA).
Task Datasets Features Timesteps Metrics Length
Imputation ETTh1/ETTh2 7 17,420 MSE, MAE 96
ETTm1/ETTm2 7 69,680
Electricity 321 26,304
Weather 21 52,696
Forecasting ETTh1/ETTh2 7 17,420 MSE, MAE 96–720
ETTm1/ETTm2 7 69,680
Electricity 321 26,304
Traffic 862 17,544
Weather 21 52,696
ILI 7 966
Classification EthanolConcentration 3 1,751 Accuracy 29-1,751
FaceDetection 144 62
Handwriting 3 152
Heartbeat 61 405
JapaneseVowels 12 29
PEMS-SF 963 144
SelfRegulationSCP1 6 896
SelfRegulationSCP2 7 1,152
SpokenArabicDigits 13 93
UWaveGestureLibrary 3 315
(b) Imputation task results. Random masking ratios of 12.5%, 25%, 37.5%, and 50% are applied. Red and blue denote first and second best results, respectively.
(Ours) (Ours) (2023) (2023) (2023) (2022) (2023) (2022)
Mask R. MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
ETTh2
ETTm1
ETTm2
ECL
Weather
We evaluate the proposed method across three fundamental time series tasks: forecasting, missing value imputation, and classification using established public benchmark datasets. Table 1 provides a detailed overview of all datasets. Missing value imputation is performed under a Missing Completely at Random (MCAR) setting, with masking ratios ranging from 12.5%, 25%, 37.5% and 50%. Benchmarks include ETTh1, ETTh2, ETTm1, ETTm2, Electricity, and Weather datasets [zhou2021informer, electricityloaddiagrams20112014_321, Kolle2025_BGCJenaWeather]. Imputation results are quantified using Mean Squared Error (MSE) and Mean Absolute Error (MAE) over sequence lengths of 96 time steps.
Long-term forecasting performance is assessed on nine widely-used benchmarks: ETTh1, ETTh2 (Electricity Transformer Temperature datasets of the China State Grid, recorded at an hourly resolution), ETTm1, ETTm2 (15-minute variants of the same transformer monitoring system), Electricity (ECL - hourly household consumption data from 321 U.S. clients), Traffic (hourly road occupancy rates from CalTrans loop detectors in the United States), Weather (10-minute meteorological observations from 21 U.S. stations), Exchange (daily currency exchange rates across major global economies), and ILI (weekly influenza-like illness records from the U.S. CDC). These datasets are extensively used for time-series benchmarking and are publicly available [wu2022autoformer]. Forecasting horizons are set between 96, 192, 336, and 720 time steps. Long-term forecasting performance is evaluated using MSE and MAE.
The multivariate time series classification is evaluated on ten diverse datasets from the University of East Anglia (UEA) archive [bagnall2018ueamultivariatetimeseries], which includes applications from healthcare to activity recognition. To rigorously assess cross-domain transferability, we utilize seven publicly available international datasets spanning diverse applications: Human Safety/Healthcare (Fall Event [BOUBEZOUL2020105577], Risky Driving [DARUS33012023]); Human Activity Recognition (HAR) (UCI-HAR [har240], PAMAP2 [pamap2_physical_activity_monitoring_231]); Agricultural Technology/Welfare (ActBe-Calf [DISSANAYAKE2025111462]); and Industrial Monitoring/Predictive Maintenance (MetroPT-3 [metropt3], NASA Turbofan Engine [NasaTurboFan]).
2.2 Imputation
Imputation is the task of filling in missing values in a dataset. In real-world time series data, sensors can fail or data can be corrupted, which may create gaps in the data. This task tests the ability of the model to intelligently reconstruct missing data based on the surrounding context, which is vital for maintaining data quality.
@@ 2.3 Long-Term Forecasting
Table 2: Multivariate long-term forecasting results. Input lookback window $L=96$ ($L=36$ for ILI). Prediction horizons $H\in\{96,192,336,720\}$ ($H\in\{24,36,48,60\}$ for ILI). Avg is averaged from all four prediction lengths. Red/Blue: First/Second ranks.
@@
Time series forecasting plays a crucial role in applications such as weather prediction, traffic management, and energy consumption planning. Long-term forecasting aims to predict future values far ahead in time. For example, it involves using historical weather data to estimate not only the temperature of tomorrow but also the conditions over the coming days or even weeks. We evaluate long-term forecasting performance using established protocols from the prior literature  [nie2023timeseriesworth64, wu2023timesnettemporal2dvariationmodeling, wu2022autoformer] across nine diverse benchmark datasets. The prediction horizons are set to $H\in\{96,192,336,720\}$ for the standard benchmarks and $H\in\{24,36,48,60\}$ for the ILI data set. As presented in Table 2, MSTN-Transformer and MSTN-BiLSTM show improved performance compared to specialized time series forecasting architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].